OpenStudy (rissyroo118):

Which is true about the line whose equation is x = 5? The slope is zero. The y-intercept is 4. The slope is undefined. The value of x always equals the value of y.

OpenStudy (ipwnbunnies):

x=constant is a vertical line through that x-value, which has as undefined slope.

OpenStudy (rissyroo118):

c again @iPwnBunnies

OpenStudy (ipwnbunnies):

Most likely.

OpenStudy (rissyroo118):

@iPwnbunnies Write the equation in standard form: 3x – 1 = 2y 3x + 2y = 1 3x – 2y = 1 3x + 2y = –1 3x – 2y = –1

OpenStudy (ipwnbunnies):

An equation in standard form occurs when all the variables are on one side, and the constants are on the other side.

OpenStudy (rissyroo118):

i got A 3x + 2y=1? @iPwnBunnies

OpenStudy (ipwnbunnies):

Almost, try again. You would have to subtract 2x from both sides and add 1 to both sides.

OpenStudy (ipwnbunnies):

subtract 2y I mean

OpenStudy (rissyroo118):

B 3x-2y=1 @iPwnBunnies

OpenStudy (ipwnbunnies):

If I was grading this, I would grade this question correct.

OpenStudy (rissyroo118):

yay hopefully i am corrct!

OpenStudy (rissyroo118):

What are the x- and y-intercepts of the graph of 6x – 5y = –30? x-int: 5, y-int: –6 x-int: –5, y-int: 6 x-int: –6, y-int: 5 x-int: 6, y-int: –5 @iPwnBunnies OK so , so i got D x-int: 6,-int: -5 im just tryna make sure i am right before plugging my answer in!

OpenStudy (ipwnbunnies):

Nope, not quite. Remember, finding x-intercept, set y equal to 0. When finding y intercept, set x equal to 0.

OpenStudy (rissyroo118):

okay so D is out of the question. i don't believe x-int: -6 y-int:5 would make much sense so either a or b?

OpenStudy (ipwnbunnies):

Yes, it's either A or B. Don't guess, do the work. It's easy. Two step problem.

OpenStudy (rissyroo118):

it's b x-int: –5, y-int: 6

OpenStudy (rissyroo118):

@iPwnBunnies

OpenStudy (ipwnbunnies):

If this was a burger, I would say you had a hard time finishing the burger, but it was worth the pain.

OpenStudy (rissyroo118):

wait am i wrong?

OpenStudy (rissyroo118):

@iPwnBunnies

OpenStudy (ipwnbunnies):

No lol. I was just saying you had a hard time with it, but came through at the end.

OpenStudy (rissyroo118):

ohhhhhh lol i was so lost when you said that! but thank you so much for your time!

OpenStudy (ipwnbunnies):

no prob